OpenStudy (anonymous):

In horse racing, a trifecta is a bet for which you must pick the first, second, and third place winners in their proper order in order to win a payoff. If there are fourteen horses racing, and you randomly select three as your bet in the trifecta, what is the probability that you will win?

OpenStudy (anonymous):

number of ways they can come in first, second and third is \(14\times 13\times12\) you have one of those so your probability is \[\frac{1}{14\times 13\times 12}\]